OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin(3x)-cos x

OpenStudy (anonymous):

sin(3x)= sin(2x+x) =sin(2x)cos(x)+sin(x)cos(2x) = [2sin(x)cos(x)]*cos(x) + {sin(x) [2cos^2(x)-1]} =2sin(x) *cos^2(x)+ 2sin(x)*cos^2(x)-sin(x) =4sin(x)cos^2(x)-sin(x)-cos(x) ,the last cos(x) from the question I hope I did not get anything wrong while typing, but anyways the idea still the same

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you are welcome :)